# LHAASO Protons versus LHAASO Diffuse Gamma Rays: A Consistency Check

### v1

@@ Abstract
We perform the first direct consistency check between the recently measured proton spectrum at the knee by LHAASO and the collaboration’s own high-precision mapping of Galactic diffuse gamma-ray emission. By modeling the hadronic gamma-ray production using the updated cosmic-ray spectra, gas templates and cross-section models, we show that the predicted gamma-ray flux robustly overshoots the LHAASO data in both inner and lateral Galactic regions. This persistent mismatch in both normalization and spectral shape challenges conventional scenarios linking the local cosmic-ray sea to Galactic gamma-ray emission, and calls for a revision of current cosmic ray models in the TeV-PeV sky.
I Introduction
@@
Recently, the Large High Altitude Air Shower Observatory (LHAASO) measured the cosmic-ray (CR) proton spectrum with unprecedented accuracy in the energy range known as the “knee” [1]. A few months ago, the same collaboration also provided an updated measurement of high-energy (TeV-PeV) Galactic diffuse gamma-ray emission [2, 3] using in combination the Water Cherenkov Detector Array (WCDA) and the Square Kilometer Array (KM2A). This diffuse gamma-ray emission is primarily generated by the interaction of cosmic rays (mainly protons and helium) with the gas present in our Galaxy.
@@
The location of the knee, around a few PeV, has long been associated with the maximum energy achievable by Galactic CR accelerators, and its precise shape carries crucial information about source spectra, propagation effects, and possible contributions from different nuclear species. LHAASO’s new proton spectrum, with its reduced statistical and systematic uncertainties, offers an opportunity to revisit existing models of CR propagation at the end of the Galactic spectrum. At the same time, the diffuse gamma-ray map measured by LHAASO traces the product of the local CR intensity and the interstellar gas distribution, providing a complementary probe of the same underlying particle population. In particular, any mismatch between the inferred CR proton flux and the secondary gamma-ray emission would point toward revisions in our understanding of, e.g., diffusion coefficients [4, 5], different gas densities along the Galactic plane, contributions from unresolved source populations [6, 7, 8, 9, 10], etc.
@@
Several theoretical studies have modeled the TeV-PeV diffuse gamma-ray flux by assuming a smooth CR sea normalized to locally measured spectra [11, 12, 13, 14], or by exploring how CR transport models, incorporating interactions with interstellar turbulence, affect this observable [11, 12, 15, 16]. Until now, however, such comparisons have been limited by either the accuracy of the proton measurements or the angular and energy resolution of the gamma-ray observations. LHAASO’s simultaneous, high-precision observations remove many of these previous ambiguities. By convolving the measured proton spectrum with a standard -decay emissivity model and folding in the latest gas survey data, one can generate a “predicted” diffuse gamma-ray intensity that is directly comparable to the LHAASO gamma-ray dataset.
@@
II The cosmic-ray spectrum at the ’knee’
Motivated by the recent discovery of rigidity‐dependent spectral breaks in nearly all CR species by direct missions (PAMELA [17], AMS-02 [18, 19], DAMPE [20, 21], CALET [22, 23], CREAM [24] and ATIC-2 [25]) we adopt the parameterization introduced in [26] to describe the proton and helium fluxes as:
@@ II The cosmic-ray spectrum at the ’knee’
As it is discussed, e.g., in [27], CR proton data by the aforementioned direct detection experiments reveal two spectral breaks at GeV and TeV. To properly account for diffuse ray emission measured by LHAASO, however, it is necessary to extend the analysis to higher proton energies. Indeed, the most relevant energy range for CR protons contributing to ray diffuse emission measured by LHAASO spans from to GeV. This follows from the fact that a photon with energy is most probably produced by a nucleon with energy in CR hadronic interactions [28].
Within this energy range, highlighted by shaded bands in Fig. 1, CR proton data show the existence of two additional spectral breaks. The first one at TeV was firstly observed by the indirect experiment GRAPES-3 [29] and it is now more precisely constrained by the results reported by the LHAASO collaboration. The second corresponds to the cosmic ray proton knee, located at around PeV, finally probed by LHAASO.
Considering the above discussion, we fit CR proton data below GeV by using a functional form featuring four spectral breaks. The resulting fit is shown with a black solid line in the upper panel of Fig. 1 while the best fit parameter values are reported in Tab. 1. Our fit includes the LHAASO measurements in the energy range from to PeV (we consider the measurements obtained with the EPOS-LHC hadronic model). Notably, these observations are in good agreement with the previous IceTop data [30] above the knee, and statistically dominate the determination of the proton spectrum at PeV energies.
@@
Best-fit values of cosmic ray fluxes | ||
Parameter | Proton | Helium |
[GeV-1m-2s-1sr-1] | ||
[TeV] | ||
[TeV] | ||
[PeV] | ||
[PeV] |
The CR helium flux is determined by following a similar procedure, i.e., we fit the available data for this component by using the functional form given in Eq. 1. Our results are shown with a black line in the lower panel of Fig. 1 alongside observational data. The best-fit parameters are listed in Tab. 1. The CR helium flux is comparable to the proton flux at low energies. However, it becomes dominant above GeV, due to the fact that the helium knee is located at a larger energy than the proton knee.
@@
This contribution can be estimated from observations of the all particle flux measured by LHAASO [32] and other detectors (HAWC [33], Tunka [34], KASCADE [35, 31] and IceTop [30]). In particular, we can write , allowing us to conclude that at low energies directly from observational data. Finally, we note that the assumption of a common rigidity spectrum naturally predicts that heavier elements become increasingly significant at higher energies since the location of their spectral knees is shifted in energy proportionally to their mass number.
@@
where is the energy per nucleon.
@@
III The Galactic diffuse gamma-ray emission
The diffuse -ray emission resulting from CR interactions with the interstellar medium can be expressed as [36, 12]:
@@ III The Galactic diffuse gamma-ray emission
The function depends solely on the photon energy and is computed using the AAFRAG parameterization [37, 38] for the photon production cross section, which we find to be in closest agreement with the approach of [39], based entirely on accelerator data fits. A comparison with alternative cross-section models is provided in the discussion section.
@@
Directional dependence, by contrast, is mainly determined by the distribution of gas and CRs in the Galaxy. The interstellar gas is mainly composed of atomic (H) and molecular hydrogen (H2), whose distributions are traced by the 21-cm [40] and CO [41] emission lines. We include both components in our analysis by the map provided by the GALPROP code111galprop.stanford.edu [42]. Alternatively, the gas distribution can be inferred from the dust opacity (, obtained at 353 GHz), measured by Planck Collaboration222esa.int/ScienceExploration/SpaceScience/Planck [43]. This quantity can be used as a tracer for the hydrogen gas column density () since dust is uniformly mixed with neutral gas. The dust-to-gas conversion factor is calibrated on experimental data; for our analysis, we use , as reported by [44]. To account for heavier elements, we scale the hydrogen density by a factor of 1.42 in both models, reflecting the Solar System composition, assumed to be representative of the entire Galactic Disk [45].
Finally, the spatial profile reflects the distribution of CRs in the Galaxy. Its specific form is determined by the spatial distribution of high energy CR sources and by CR propagation in the galactic magnetic field, see e.g., [12] for approximate expressions. For simplicity, in this work we initially assume that CR are uniformly distributed in the Galaxy (i.e. ) and later address how potential deviations from this assumption could impact our results. Under this assumption, the function essentially describes the gas column density along a given line of sight. It is important to note that any deviation from would alter only the angular structure of the diffuse gamma-ray flux, not its energy spectrum. Therefore, spectral discrepancies between the predicted and observed gamma-ray fluxes can serve as indicators of a position-dependent CR spectrum in the Galaxy.
IV Results
The diffuse -ray fluxes computed using the methodology discussed in the previous section are presented in Fig. 3 for both sky regions probed by LHAASO and after applying the same mask used by the collaboration in Ref. [3]. Two shaded bands, which represent our uncertainty in CR composition and contribution of heavy elements, illustrate our predictions based on different assumptions for the interstellar gas distributions. The blue band is obtained by using the gas distribution provided with the GALPROP code. The green band is based instead on the gas column density inferred from the Planck dust opacity () map.
@@ IV Results
It is important to note that the observation of an excess of -rays, as recently discussed in [14] and in the original LHAASO -ray papers [2, 3], may suggest the presence of a missing diffuse -ray component or a non-standard cosmic ray diffusion paradigm. However, the evidence of a deficit in observed emission represents a new and more puzzling feature.
We find this result cannot be attributed to the hadronic interaction model adopted for the diffuse flux calculations. In order to show this, we compute the diffuse -flux expectation using other cross-section parametrizations in the literature. Specifically, we consider Geant4, Pythia8, SIBYLL and QGSJET as defined in [46]) and we include the results in the Appendix. We see that the AAFrag cross-section gives the smaller (and softer) -ray predictions and, therefore, the tension with LHAASO -ray observations would increase for other parameterizations.
@@
Finally, the observation of a energy-dependent flux ratio between the two sky regions could be a hint for a spatial dependence of the CR spectral index in the Galaxy. Previous works ([15, 47, 48, 49, 36, 50, 51, 11, 12]) have explored this possibility, based on the fact that Fermi-LAT ray data at GeV energies suggest CR spectral hardening in the direction of the Galactic center. We note, however, that diffuse -ray observations by LHAASO do not distinguish between these spatial-dependent CR transport models and the conventional scenario (as was already noted in [14]), being well compatible with a constant within uncertainties.
V Discussion
@@
We have obtained an over-prediction of the -flux compared to data which is particularly evident in the lateral LHAASO observation region and when using GALPROP gas distribution. We highlight that this tension is difficult to alleviate. An overall normalization difference can only be attributed to the uncertainty on the Galactic gas or variations of the CR density in the Galaxy. One possibility that has been discussed in [12, 14] is that the cosmic ray density is not uniform in the Galaxy but resembles the distribution of sources (mainly supernova remnants, SNR), motivated by a possible confinement in the proximity of the sources. However, including this variation in our analysis would cause a reduction of the gamma ray flux of up to 20 in the lateral region (in the most optimistic scenarios) while it would increase the prediction in the inner region, worsening the tension.
@@
On the other hand, the discrepancy of the predicted spectral shape can be related to uncertainties on hadronic interaction models, -ray absorption, spatial dependence of the CR energy spectrum or the contribution of an additional diffuse component. Nevertheless, we have shown that the adoption of other cross-sections would increase the -flux while predicting a slightly harder slope. Similarly, an additional diffuse component contributing to the LHAASO measurement would require a lower -ray flux produced by CRs. Therefore, both options would enlarge the discrepancy found in this work.
@@
Two alternatives that were not fully explored in this paper are the additional -ray absorption by interstellar radiation field (ISRF) and the possibility of a spatial-dependent CR energy spectrum. Absorption of -rays with energies TeV, due to kinematical threshold for pair production, would require a population of background photons more energetic than CMB. Their contribution is, however, believed to be small according to present estimates, see e.g.[52].
@@
To assess the possibility of CR-spectrum spatial dependency, we have computed the ratio between the diffuse -ray emission measured by LHAASO in the two different sky regions. We note due to the large uncertainties, LHAASO data do not distinguish a hint of spatial dependent spectral indices in the TeV-PeV energy range. The ratio among the current data set from the two regions, as reported in Fig.5, is consistent with a constant value, which indicates that, in principle, the cosmic ray (CR) spectrum in both regions is the same.
@@
Conversely, the ratios between the models and the data shown in Fig.4 reveal a significant mismatch between the spectral shape of local CR and the measured gamma rays. One possible explanation is that the CR spectrum we are measuring may not be the one responsible for the observed gamma-ray emission. Specifically, the CR spectrum in the remainder of the Galaxy could be characterized by a different knee location at around TeV, which is compatible with the observed gamma break. This idea, partially explored in previous work [53], could be directly tested in the future through gamma-ray observations.
@@
The uncertainty in the proton flux can potentially affect our conclusions. Previous measurements by KASCADE [35, 31] reported a significantly lower proton flux than IceTop in the PeV energy range. A recent work in Ref. [16] highlighted that the flux measured by LHAASO-KM2A was incompatible with the IceTop proton flux dataset while in good agreement with the KASCADE data. Notably, a recent reanalysis of the KASCADE proton flux [54] finds consistency with both IceTop and LHAASO measurements. If this updated dataset is confirmed by the KASCADE collaboration, then all charged cosmic-ray results would be mutually compatible, making the persistent tension with the diffuse gamma-ray observations even more puzzling.
Conversely, the helium flux lacks measurements between 100 TeV and a few PeV, providing another potential source of uncertainty in all models so far. Measurements of the cosmic ray composition (given by ) by LHAASO together with their proton flux seem to suggest that the cosmic ray flux is dominated by protons between 1-10 PeV. These measurements do not exclude the possibility of a helium contribution at the knee lower than modeled in this work, potentially decreasing the diffuse -ray flux. Dedicated helium flux measurements in the 100 TeV–PeV band are therefore essential [55]; once available, they will be incorporated into our framework in future work.
Appendix
Computed diffuse -ray fluxes for multiple cross-section parametrization in inner and lateral LHAASO regions displayed in Fig. 6. Maximum -ray prediction is found to come from SIBYLL parametrization while AAFrag gives the minimal prediction at -ray energies TeV.

### v2

# LHAASO protons vs LHAASO diffuse gamma-rays

@@ Abstract
We perform the first direct consistency check between the recently measured proton spectrum at the knee by Large High Altitude Air Shower Observatory (LHAASO) and the collaboration’s own high-precision mapping of Galactic diffuse gamma-ray emission. By modeling the hadronic gamma-ray production using the updated cosmic-ray spectra, gas templates and cross-section models, we show that the predicted gamma-ray flux robustly overshoots the LHAASO data in both inner and lateral Galactic regions. This persistent mismatch in both normalization and spectral shape challenges conventional scenarios linking the local cosmic-ray sea to Galactic gamma-ray emission, and calls for a revision of current cosmic ray models in the TeV-PeV sky.
keywords:
astroparticle physics – cosmic rays – gamma-rays: diffuse background1 Introduction
@@
Recently, the Large High Altitude Air Shower Observatory (LHAASO) measured the cosmic-ray (CR) proton spectrum with unprecedented accuracy in the energy range known as the ‘knee’ (Cao et al., 2025a). A few months ago, the same collaboration also provided an updated measurement of high-energy (TeV-PeV) Galactic diffuse gamma-ray emission (Cao et al., 2023, 2025b) using in combination the Water Cherenkov Detector Array (WCDA) and the Square Kilometer Array (KM2A). This diffuse gamma-ray emission is primarily generated by the interaction of cosmic rays (mainly protons and helium) with the gas present in our Galaxy.
@@
The location of the knee, around a few PeV, has long been associated with the maximum energy achievable by Galactic CR accelerators, and its precise shape carries crucial information about source spectra, propagation effects, and possible contributions from different nuclear species. LHAASO’s new proton spectrum, with its reduced statistical and systematic uncertainties, offers an opportunity to revisit existing models of CR propagation at the end of the Galactic spectrum. At the same time, the diffuse gamma-ray map measured by LHAASO traces the product of the local CR intensity and the interstellar gas distribution, providing a complementary probe of the same underlying particle population. In particular, any mismatch between the inferred CR proton flux and the secondary gamma-ray emission would point toward revisions in our understanding of, e.g., diffusion coefficients (Ptuskin et al., 1993; Syrovatskii, 1971), different gas densities along the Galactic Plane, contributions from unresolved source populations (Steppa & Egberts, 2020; Cataldo et al., 2020; Vecchiotti et al., 2022; Lipari & Vernetto, 2025; He et al., 2025), etc.
@@
Several theoretical studies have modeled the TeV-PeV diffuse gamma-ray flux by assuming a smooth CR sea normalized to locally measured spectra (Lipari & Vernetto, 2018; Cataldo et al., 2019; Schwefer et al., 2023; Vecchiotti et al., 2024), or by exploring how CR transport models, incorporating interactions with interstellar turbulence, affect this observable (Lipari & Vernetto, 2018; Cataldo et al., 2019; Gaggero et al., 2015; Luque et al., 2025). Until now, however, such comparisons have been limited by either the accuracy of the proton measurements or the angular and energy resolution of the gamma-ray observations. LHAASO’s simultaneous, high-precision observations remove many of these previous ambiguities. By convolving the measured proton spectrum with a standard -decay emissivity model and folding in the latest gas survey data, one can generate a “predicted” diffuse gamma-ray intensity that is directly comparable to the LHAASO gamma-ray dataset.
@@
2 The cosmic-ray spectrum at the ’knee’
Motivated by the recent discovery of rigidity-dependent spectral breaks in nearly all CR species by direct missions (PAMELA Adriani et al. 2011, AMS-02 Aguilar et al. 2015a, b, DAMPE An et al. 2019; Alemanno et al. 2021, CALET Adriani et al. 2019, 2023, CREAM Yoon et al. 2017 and ATIC-2 Panov et al. 2009) we adopt the parameterization introduced in Ter-Antonyan & Haroyan (2000) to describe the proton and helium fluxes as:
@@ 2 The cosmic-ray spectrum at the ’knee’
As it is discussed, e.g., in Lipari & Vernetto (2020), CR proton data by the aforementioned direct detection experiments reveal two spectral breaks at GeV and TeV. To properly account for diffuse ray emission measured by LHAASO, however, it is necessary to extend the analysis to higher proton energies. Indeed, the most relevant energy range for CR protons contributing to ray diffuse emission measured by LHAASO spans from to GeV. This follows from the fact that a photon with energy is most probably produced by a nucleon with energy in CR hadronic interactions (Kelner et al., 2006).
Within this energy range, highlighted by shaded bands in Fig. 1, CR proton data show the existence of two additional spectral breaks. The first one at TeV was firstly observed by the indirect experiment GRAPES-3 (Varsi et al., 2024) and it is now more precisely constrained by the results reported by the LHAASO collaboration. The second corresponds to the cosmic ray proton knee, located at around PeV, finally probed by LHAASO.
Considering the above discussion, we fit CR proton data below GeV by using a functional form featuring four spectral breaks. The resulting fit is shown with a black solid line in Fig. 1 while the best fit parameter values are reported in the Supplementary Material. Our fit includes the LHAASO measurements in the energy range from to PeV (we consider the measurements obtained with the EPOS-LHC hadronic model). Notably, these observations are in good agreement with the previous IceTop data (Aartsen et al., 2019) above the knee, and statistically dominate the determination of the proton spectrum at PeV energies.
@@
The CR helium flux is determined by following a similar procedure, i.e., we fit the available data for this component by using the functional form given in Eq. 1. Our results are shown in the Supplementary Material, alongside the best-fit parameters. The CR helium flux is comparable to the proton flux at low energies. However, it becomes dominant above GeV, due to the fact that the helium knee is located at a larger energy than the proton knee.
@@
This contribution can be estimated from observations of the all particle flux measured by LHAASO (Cao et al., 2024) and other detectors (HAWC Morales-Soto et al. (2021), Tunka Prosin et al. (2014), KASCADE Antoni et al. (2005); Finger (2011) and IceTop Aartsen et al. (2019)). In particular, we can write , allowing us to conclude that at low energies directly from observational data. Finally, we note that the assumption of a common rigidity spectrum naturally predicts that heavier elements become increasingly significant at higher energies since the location of their spectral knees is shifted in energy proportionally to their mass number.
@@
where is the energy per nucleon. The factor comes from considering that the number of nucleons () in a nucleus of a atomic mass is , while the energy per nucleon is . Therefore, the differential flux of nucleons scales as .
@@
3 The Galactic diffuse gamma-ray emission
The diffuse -ray emission resulting from CR interactions with the interstellar medium can be expressed as in Pagliaroli et al. (2016); Cataldo et al. (2019):
@@ 3 The Galactic diffuse gamma-ray emission
The function depends solely on the photon energy and is computed using the AAFRAG parameterization (Koldobskiy et al., 2021; Kachelriess et al., 2023) for the photon production cross section, which we find to be in closest agreement with the approach of Orusa et al. (2023), based entirely on accelerator data fits. A comparison with alternative cross-section models is provided in the discussion section.
@@
Directional dependence, by contrast, is mainly determined by the distribution of gas and CRs in the Galaxy. The interstellar gas is mainly composed of atomic (H) and molecular hydrogen (H2), whose distributions are traced by the 21-cm (Ben Bekhti et al., 2016) and CO (Dame et al., 2001) emission lines. We include both components in our analysis by the map provided by the GALPROP code111galprop.stanford.edu (Porter et al., 2022). Alternatively, the gas distribution can be inferred from the dust opacity (, obtained at 353 GHz), measured by Planck Collaboration222esa.int/ScienceExploration/SpaceScience/Planck (Aghanim et al., 2016). This quantity can be used as a tracer for the hydrogen gas column density () since dust is uniformly mixed with neutral gas. The dust-to-gas conversion factor is calibrated on experimental data; for our analysis, we use , as reported by Ade et al. (2011). To account for heavier elements, we scale the hydrogen density by a factor of 1.42 in both models, reflecting the Solar System composition, assumed to be representative of the entire Galactic Disk (Ferriere, 2001).
Finally, the spatial profile reflects the distribution of CRs in the Galaxy. Its specific form is determined by the spatial distribution of high energy CR sources and by CR propagation in the Galactic magnetic field, see e.g., Cataldo et al. (2019) for approximate expressions. For simplicity, in this work we initially assume that CR are uniformly distributed in the Galaxy (i.e. ) and later address how potential deviations from this assumption could impact our results. Under this assumption, the function essentially describes the gas column density along a given line of sight. It is important to note that any deviation from would alter only the angular structure of the diffuse gamma-ray flux, not its energy spectrum. Therefore, spectral discrepancies between the predicted and observed gamma-ray fluxes can serve as indicators of a position-dependent CR spectrum in the Galaxy.
4 Results
The diffuse -ray fluxes computed using the methodology discussed in the previous section are presented in Fig. 3 for both sky regions probed by LHAASO and after applying the same mask used by the collaboration in Cao et al. (2025b). Two shaded bands, which represent our uncertainty in CR composition and contribution of heavy elements, illustrate our predictions based on different assumptions for the interstellar gas distributions. The blue band is obtained by using the gas distribution provided with the GALPROP code. The green band is based instead on the gas column density inferred from the Planck dust opacity () map.
@@ 4 Results
It is important to note that the observation of an excess of -rays, as recently discussed in Vecchiotti et al. (2024) and in the original LHAASO -ray papers (Cao et al., 2023, 2025b), may suggest the presence of a missing diffuse -ray component or a non-standard cosmic ray diffusion paradigm. However, the evidence of a deficit in observed emission represents a new and more puzzling feature.
We find this result cannot be attributed to the hadronic interaction model adopted for the diffuse flux calculations. In order to show this, we compute the diffuse -flux expectation using other cross-section parametrizations in the literature. Specifically, we consider Geant4, Pythia8, SIBYLL and QGSJET as defined in Kafexhiu et al. (2014)) and we include the results in the Appendix. We see that the AAFrag cross-section gives the smaller (and softer) -ray predictions and, therefore, the tension with LHAASO -ray observations would increase for other parameterizations.
@@
Finally, the observation of a energy-dependent flux ratio between the two sky regions could be a hint for a spatial dependence of the CR spectral index in the Galaxy. Previous works (Gaggero et al., 2015; Acero et al., 2016; Yang et al., 2016; Pothast et al., 2018; Pagliaroli et al., 2016; Recchia et al., 2016; Cerri et al., 2017; Lipari & Vernetto, 2018; Cataldo et al., 2019) have explored this possibility, based on the fact that Fermi-LAT ray data at GeV energies suggest CR spectral hardening in the direction of the Galactic Center. We note, however, that diffuse -ray observations by LHAASO do not distinguish between these spatial-dependent CR transport models and the conventional scenario (as was already noted in Vecchiotti et al. (2024)), being well compatible with a constant within uncertainties.
5 Discussion
@@
We have obtained an over-prediction of the -flux compared to data which is particularly evident in the lateral LHAASO observation region and when using GALPROP gas distribution. We highlight that this tension is difficult to alleviate. An overall normalization difference can only be attributed to the uncertainty on the Galactic gas or variations of the CR density in the Galaxy. One possibility that has been discussed in Cataldo et al. (2019); Vecchiotti et al. (2024) is that the cosmic ray density is not uniform in the Galaxy but resembles the distribution of sources (mainly supernova remnants, SNR), motivated by a possible confinement in the proximity of the sources. However, including this variation in our analysis would cause a reduction of the gamma ray flux of up to 20 in the lateral region (in the most optimistic scenarios) while it would increase the prediction in the inner region, worsening the tension.
@@
On the other hand, the discrepancy of the predicted spectral shape can be related to uncertainties on hadronic interaction models, -ray absorption, spatial dependence of the CR energy spectrum or the contribution of an additional diffuse component (e.g. a population of unresolved sources). Nevertheless, we have shown that the adoption of other cross-sections would increase the -flux while predicting a slightly harder slope. Similarly, an additional diffuse component contributing to the LHAASO measurement would require a lower -ray flux produced by CRs, especially in the inner region where a higher density of unresolved sources could be expected. Therefore, both options would enlarge the discrepancy found in this work.
@@
Two alternatives that were not fully explored in this paper are the additional -ray absorption by interstellar radiation field (ISRF) and the possibility of a spatial-dependent CR energy spectrum. Absorption of -rays with energies TeV, due to kinematical threshold for pair production, would require a population of background photons more energetic than CMB. Their contribution is, however, believed to be small according to present estimates, see e.g.Vernetto & Lipari (2016).
@@
To assess the possibility of CR-spectrum spatial dependency, we have computed the ratio between the diffuse -ray emission measured by LHAASO in the two different sky regions. We note due to the large uncertainties, LHAASO data do not distinguish a hint of spatial dependent spectral indices in the TeV-PeV energy range. The ratio among the current data set from the two regions, as reported in Fig.5, is consistent with a constant value, which indicates that, in principle, the cosmic ray (CR) spectrum in both regions is the same. Nevertheless, the ratio predicted from both Galactic gas maps is considerably below the ratio obtained by observations at gamma-ray energies TeV. This difference in normalization could be attributed to a slightly higher density of gas or CRs in the inner region. However, this would increase the high-energy gamma-ray flux ( TeV) in the inner region, possibly leading to a predicted flux above the LHAASO observations.
Conversely, the ratios between the models and the data shown in Fig.4 reveal a significant mismatch between the spectral shape of local CR and the measured gamma rays. This discrepancy is most noticeable in the inner region for gamma-ray energies TeV, while is the lowest in the lateral region.
@@
One possible explanation is that the CR spectrum we are measuring may not be the one responsible for the observed gamma-ray emission. Specifically, the CR spectrum in the remainder of the Galaxy could be characterized by a different knee location at around TeV, which is compatible with the observed gamma break. This idea, partially explored in previous work (Prévotat et al., 2024), could be directly tested in the future through gamma-ray observations.
@@
The uncertainty in the proton flux can potentially affect our conclusions. Previous measurements by KASCADE (Antoni et al., 2005; Finger, 2011) reported a significantly lower proton flux than IceTop in the PeV energy range. A recent work in Ref. Luque et al. (2025) highlighted that the flux measured by LHAASO-KM2A was incompatible with the IceTop proton flux dataset while in good agreement with the KASCADE data. Notably, a recent reanalysis of the KASCADE proton flux (Kuznetsov et al., 2024) finds consistency with both IceTop and LHAASO measurements. If this updated dataset is confirmed by the KASCADE collaboration, then all charged cosmic-ray results would be mutually compatible, making the persistent tension with the diffuse gamma-ray observations even more puzzling.
Conversely, the helium flux lacks measurements between 100 TeV and a few PeV, providing another potential source of uncertainty in all models so far. Measurements of the cosmic ray composition (given by ) by LHAASO together with their proton flux seem to suggest that the cosmic ray flux is dominated by protons between 1-10 PeV. These measurements do not exclude the possibility of a helium contribution at the knee lower than modeled in this work, potentially decreasing the diffuse -ray flux. Dedicated helium flux measurements in the 100 TeV–PeV band are therefore essential (Wang et al., 2024); once available, they will be incorporated into our framework in future work.
Acknowledgements
The Authors thank Prof. Yuhui Li for providing the LHAASO mask needed to correctly perform the estimation of the diffuse emission. The work of GP and FLV is partially supported by grant number 2022E2J4RK "PANTHEON: Perspectives in Astroparticle and Neutrino THEory with Old and New messengers" under the program PRIN 2022 funded by the Italian Ministero dell’Università e della Ricerca (MUR) and by the European Union – Next Generation EU.
Data Availability
All external data used in this work were obtained from the cited sources. Data produced in the analysis of this study are available upon reasonable request.